OpenStudy (anonymous):

@callisto, Help! How do i find the exact value of: tan(135°) + sin(30°) cot (120°):

OpenStudy (anonymous):

scientific caclulator. then do proecesses to decimals

OpenStudy (anonymous):

put value for sin 30 and take cot as cot(60+60)

OpenStudy (callisto):

tan(135°) = -tan 45 =-1 sin(30°) =1/2 cot (120°) = 1/tan120 = 1/ -tan60 = 1/-sqrt 3 Can you work this out?

OpenStudy (anonymous):

k, one sec

OpenStudy (anonymous):

tan is added with the other values, is sin multiplied by cot?

OpenStudy (lgbasallote):

haizzz wouldn't it be nice to be personally tagged in a question *sigh* haha

OpenStudy (callisto):

yes

OpenStudy (callisto):

@Open2study

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

am i solving by decimal approximation?

OpenStudy (callisto):

Hmmm... tan(135°) + sin(30°) cot (120°): = -tan(180 - 135) + sin30 (-cot (180-120)) = -tan45 + sin30 (-cot60) = -1 + (1/2)(-1/tan60) = -1 + (1/2) (-1/sqrt3) Understand it? and can you try to simplify the answer?

OpenStudy (anonymous):

Is it sqrt (3)/2 - 1?

OpenStudy (callisto):

Not really... Can you try again??

OpenStudy (anonymous):

If its in decimals, is the answer -0.1339?

OpenStudy (callisto):

Nope...

OpenStudy (anonymous):

Is it 1/2 (√3 - 2)?

OpenStudy (callisto):

Just a minute!

OpenStudy (anonymous):

ok

OpenStudy (callisto):

Finally, I decided not to trust my calculator :S

OpenStudy (anonymous):

Omg, You went through so much trouble scanning and all just to help me! Thank you so much Jos!

OpenStudy (callisto):

You're welcome.....

OpenStudy (anonymous):

one question, Hoe did it go from -2√3 - 1 / 2√3 to -2√3 (√3) -√3 / 2(3)?

OpenStudy (anonymous):

so the exact value is -6 √-3 / 6

OpenStudy (callisto):

I think so... and I hope so....

OpenStudy (anonymous):

Wow. Thank you so much again @Callisto! You have helped me so much!

OpenStudy (callisto):

Welcome :)

OpenStudy (anonymous):

@callisto, Im finding the exact value of sin (30)

OpenStudy (anonymous):

isnt it sin(60) = sqrt(3)/2?

OpenStudy (anonymous):

it lies on quadrant I, positive

OpenStudy (callisto):

yess

OpenStudy (anonymous):

How come in your answer its 1/2?

OpenStudy (callisto):

sin(60) = sqrt(3)/2 <- correct BUT sin 30 = 1/2 ....

OpenStudy (anonymous):

Ohhh. ok, nvm I was concentrating too much on one thing. So sorry to bother you again

OpenStudy (callisto):

That's not bothering :P

OpenStudy (anonymous):

Thank you for the clarification. I get lost many times, and some times for no reason :P

OpenStudy (callisto):

Can you get it now?

OpenStudy (anonymous):

@callisto, sorry! this is the last time I will call for your assistance today, I am going to be now. I am just double checking before submitting. How come in the procedure, you are constantly adding sqrt (3), the ones written in red?

OpenStudy (callisto):

Do the rationalisation

OpenStudy (callisto):

my teacher said it's not good to leave a sqrt in the denominator, so it's my practice to do the rationalisation

OpenStudy (anonymous):

Oh! Thats right! Im sorry, Im just really sleepy so Im not thinking straight. Thank you so much Callisto! I have to go to bed no:) Thank you!

OpenStudy (callisto):

welcome ! Good night :)